# Fact-checking information generated by a large language model can decrease news discernment

### v3

@@ Abstract
Fact checking can be an effective strategy against misinformation1, 2, 3, but its implementation at scale is impeded by the overwhelming volume of information online4. Recent artificial intelligence (AI) language models have shown impressive ability in fact-checking tasks5, 6, but how humans interact with fact-checking information provided by these models is unclear. Here, we investigate the impact of fact-checking information generated by a popular large language model (LLM) on belief in, and sharing intent of, political news in a preregistered randomized control experiment. Although the LLM performs reasonably well in debunking false headlines, we find that it does not significantly affect participants’ ability to discern headline accuracy or share accurate news. Subsequent analysis reveals that the AI fact-checker is harmful in specific cases: it decreases beliefs in true headlines that it mislabels as false and increases beliefs in false headlines that it is unsure about. On the positive side, the AI fact-checking information increases sharing intents for correctly labeled true headlines. When participants are given the option to view LLM fact checks and choose to do so, they are significantly more likely to share both true and false news but only more likely to believe false news. Our findings highlight an important source of potential harm stemming from AI applications and underscore the critical need for policies to prevent or mitigate such unintended consequences.
@@
This study explores the effects of fact-checking information generated by large language models (LLMs) on public perception and dissemination of political news. Despite the public’s growing adoption of this artificial intelligence (AI) technology, and practitioners testing its ability to combat online misinformation, little is known about how people respond to LLM-driven fact-checking. This randomized control experiment reveals that even LLMs that effectively debunk false headlines do not necessarily enhance users’ abilities to discern news accuracy or promote accurate news sharing. Notably, LLM fact checks can inadvertently reduce trust in true news wrongly labeled as false and increase belief in dubious headlines when the AI is unsure about an article’s veracity. These findings highlight the urgent need for thorough research into the unintended impacts of AI fact-checking, guiding the development of policies that ensure this new technology bolsters information integrity in our increasingly digital world.
@@
Digital misinformation has rapidly become a critical issue of modern society7, 8. Recent work suggests that misinformation can erode support for climate change9, 10, contribute to vaccine hesitancy11, 12, 13, exacerbate political polarization14, and even undermine democracy15. As a mitigation strategy, fact checking has proved effective at reducing people’s belief in3, 1, 16 and intention to share17 misinformation in various cultural settings2. However, this approach is not scalable, greatly limiting its applications4.
@@
To tackle this challenge, researchers and social media platforms have been exploring automated methods18 to detect misinformation19, 20 and fact-check claims21, 22, 18, 23, 24, 25. A robust fact-checking system must possess the ability to detect claims, retrieve relevant evidence, assess the veracity of each claim, and yield justifications for the provided conclusions26, 27. Previous work attempting to meet these goals typically adopts cutting-edge artificial intelligence (AI) methods, specifically natural language processing. Nevertheless, building a functional system that can handle the vast volume of digital information on the internet, spanning various contexts and languages, remains a daunting task.
@@
Recent advances in large language models (LLMs) may appear to provide a feasible path forward. Trained on massive datasets of text from the internet, including news articles, books, and websites28, these models are knowledgeable about a wide range of topics and have shown impressive performance on tasks such as text summarization and named entity recognition29, 30. Outside the laboratory, LLMs have demonstrated remarkable abilities, even passing challenging exams designed for humans31, 32.
@@
Analyses of ChatGPT, a prominent LLM, suggest it can rate the credibility of news outlets33 and fact-check claims5, 34 without special configuration. Messages provided by LLMs to correct social media misinformation can be better than corrective messages generated by humans6. These models can generate convincing justifications for the information they provide and even engage in conversations with users to provide additional context and facilitate understanding in multiple languages. Such capabilities of LLMs, coupled with open-sourcing efforts35, 36, create a favorable environment for the development of scalable and reliable AI systems that can verify nearly all claims on the internet.
@@
However, realizing this potential requires humans to integrate LLMs into the digital information ecosystem effectively. Unfortunately, human-AI interaction is notoriously complex37. Prior work has shown that AI is often seen as objective38, 39, 40, 41, yet trust in AI depends on various factors such as individual expectations42, 43, system interactivity44, 45, and whether the AI provides information about its recommendations46, 47.
@@
In the present context, it remains unclear how humans would interact with fact-checking information provided by state-of-the-art LLMs. Therefore, a thorough analysis of this misinformation intervention is necessary before deploying models in the wild. To this end, we conduct a preregistered48, randomized controlled experiment to examine the causal effects of viewing fact-checking information provided by ChatGPT on individual beliefs in and intention to share political news. We selected ChatGPT for our study despite it not being specifically tailored for fact-checking. This decision was driven by its widespread public availability and use as well as the promising results emerging from tests of its claim verification capabilities at the time34, 49, 5.
@@ Results
We recruited a representative sample of $N=1{,}548$ U.S. participants (see Methods for more information). All participants were presented with the same 40 real political news stories, which included a headline, lede sentence (if present), and image. Half of these headlines were true and the other half were false. Half were favorable to Democrats and the other half were favorable to Republicans (see Methods for details).
Participants were separated into “belief” and “sharing” groups in which they were asked to indicate, respectively, whether they believed headlines to be accurate or would be willing to share them on social media. These questions were asked separately as priming individuals to think about headline veracity can alter sharing behavior50, 4. Three conditions were designed for each group, consisting of a control and two treatment conditions that aimed to simulate hypothetical scenarios for an automated fact-checking system on a social media platform: treated subjects were either forced to view fact-checking information provided by ChatGPT (forced condition) or given the option to reveal that information by clicking a button (optional condition). ChatGPT fact-checking information was identical for all treated subjects and presented directly below the corresponding headline (see Methods for more details). Subjects in the control condition were only shown headlines and asked the belief/sharing question without being exposed to any fact-checking information. Study participants were assigned randomly to one of the six experimental conditions. The experimental design is illustrated in Fig. 1a. See Methods for more details.
Unless otherwise stated, all $P$ values presented here are generated with two-tailed Mann-Whitney U tests and adjusted with Bonferroni correction for multiple comparisons. In Supplementary information, we also report on linear regression for all results, employing robust standard errors clustered on participant and headline.
@@
Although limited in size, our set of headlines provides us with a balanced representation of political biases and factual accuracy to evaluate the LLM. Overall, this analysis suggests that the LLM is an accurate fact-checker for false content. For true headlines, it is less accurate but can generally identify and explain when it cannot provide accurate fact-checking information. These results align with earlier studies that delve into the accuracy of LLM fact-checking utilizing much larger datasets49, 5, 34.
@@
To evaluate the effectiveness of a misinformation intervention, it is crucial to measure its impact on belief in and sharing of both true and false headlines51, 50. Although the veracity of headlines may not always fit neatly into true and false categories, as in the case of rumors with unclear veracity that are later clarified52, this framework defines the desired outcome: an effective misinformation intervention should enhance individuals’ ability to distinguish between true and false headlines such that they believe/share more accurate news.
@@
To capture the causal effects of LLM-generated fact-checking information, we compare the average discernment of participants in the treatment conditions (forced and optional) to those in the control condition. Discernment is defined as the difference between the proportion of true and false headlines that participants believe (or are willing to share), capturing the intervention’s impact on both news categories.
@@
Figure 1 (panels c,d) illustrates the effects of LLM fact checking on belief in and intent to share true and false headlines under each condition, including the mean group discernment as an annotation. In contrast with our preregistered expectations, discernment within both the belief and sharing groups was unaffected by the treatment, regardless of condition. In the belief group (Fig. 1c), participants who were forced to view AI fact checks displayed a slight mean reduction ($-4.50$%) in discernment when compared to the control group ($U=31{,}993$, $P=0.45$, $d=-0.15$, 95% CI: $[-10.04\%,0.90\%]$). The discernment of those given the option to view fact checks in this group was virtually unaffected, decreasing on average by only $-0.27$% ($U=31{,}265$, $P=1.00$, $d=-0.01$, 95% CI: $[-5.52\%,5.03\%]$).
@@
Similarly in the sharing group (Fig. 1d), we do not observe any significant changes in discernment for either treatment condition. Compared to the control group, the forced and optional conditions displayed a mean reduction of $-0.43$% ($U=35{,}785$, $P=1.00$, $d=-0.02$, 95% CI: $[-3.93\%,3.05\%]$) and $-0.67$% ($U=34{,}860$, $P=1.00$, $d=-0.03$, 95% CI: $[-4.18\%,2.92\%]$), respectively.
@@
In summary, these results indicate that the fact-checking information provided by the LLM is not an effective intervention against misinformation. This is unexpected considering that the AI provides participants with useful information, particularly for false headlines. However, this analysis does not account for the accuracy of the AI’s responses, nor does it examine how behaviors vary when individuals choose to view or not view this information. To delve deeper into these dynamics, we have supplemented our preregistered design with two additional exploratory analyses in the sections that follow.
@@
Here, we explore the causal effects of viewing LLM fact-checking information when accounting for model accuracy. The judgments made by ChatGPT for both true and false headlines fall into one of three categories: correct, incorrect, or unsure. This results in six different scenarios (True/False $\times$ Correct/Incorrect/Unsure) in which effects may be observed. However, our data contain no false headlines judged by ChatGPT to be true, resulting in five scenarios for each previously considered comparison (Belief/Sharing $\times$ Control/Optional/Forced).
To evaluate the potential impact of LLM-generated fact checks, we compare the forced and control conditions in these five scenarios, as illustrated in Fig. 2. Annotations indicate mean group differences and highlight the significant effects identified through Bonferroni-adjusted Mann-Whitney U tests.
@@
In the belief group, we found significant undesirable effects showing that LLM fact checks decreased participants’ discernment. Specifically, there was a 12.75% decrease in the belief of true headlines incorrectly judged as false by ChatGPT ($U=35{,}937$, $P<0.001$, $d=-0.38$, 95% CI: $[-18.67\%,-6.89\%]$) and a 9.12% increase in the belief of false headlines where the AI expressed uncertainty ($U=25{,}931$, $P=0.030$, $d=0.22$, 95% CI: $[1.69\%,16.35\%]$). Both cases demonstrate a behavioral change that is counter to the intended outcome of a misinformation intervention.
Regarding the sharing group, we observed mixed results. While there was an 11.09% increase in participants’ intention to share true headlines correctly judged by ChatGPT ($U=30{,}897$, $P=0.017$, $d=0.26$, 95% CI: $[4.02\%,18.06\%]$), there was also a 9.77% increase in the intention to share false headlines where ChatGPT expressed uncertainty ($U=31{,}856$, $P=0.051$, $d=0.22$, 95% CI: $[2.31\%,17.25\%]$). The former increases sharing discernment, while the latter reduces it by a similar amount.
@@
We next analyze participants’ behavior in the optional condition, comparing those who opt in to see LLM fact-checking information versus those who opt out. Figure 3 illustrates the belief in and intention to share headlines for which subjects chose to see versus not see LLM fact checks, for both true and false headlines. Each subject’s contribution to the group mean values and confidence intervals are weighted by the number of times they chose to see (or not see) each type of headline. Figure 3a shows that participants who chose to see LLM fact checks were significantly more likely to believe false headlines (29.47% increase; $U=25{,}043$, $P<0.001$, $d=0.64$, 95% CI: $[21.22\%,37.96\%]$), whereas no significant difference in belief was observed for true headlines (1.50% increase; $U=19{,}871$, $P=1.00$, $d=0.03$, 95% CI: $[-7.41\%,10.33\%]$). On the other hand, Figure 3b shows that participants who viewed LLM fact checks were more willing to share headlines from both categories, with a 34.06% increase for true headlines ($U=25{,}584$, $P<0.001$, $d=0.70$, 95% CI: $[24.68\%,43.34\%]$) and a 37.61% increase for false headlines ($U=25{,}426$, $P<0.001$, $d=0.74$, 95% CI: $[28.26\%,46.77\%]$).
Figure 3: Proportions of headlines that participants in the optional condition indicated they (a) believed or (b) were willing to share on social media, based on headline veracity and whether participants chose to see LLM fact-checking information (opt in) or not (opt out). Error bars represent 95% confidence intervals, calculated using a weighted bootstrapping method with 5,000 resamples. The mean difference between opt-in and opt-out groups (rounded to whole percentage) is annotated for each condition.
We note that this particular within-group analysis does not allow us to identify causal effects because participants are not randomly assigned to the treatment (opt in) or comparison (opt out) group for each headline. Nonetheless, when participants viewed LLM-generated fact-checking information they were much more likely to share both true and false news, but only more likely to believe false news.
@@
Our preregistered analyses also examined the potential roles played by individual attitudes toward AI and by the partisan congruence of headlines. We find evidence of both variables impacting results in certain cases when individuals are given the choice to view LLM fact-checking information. Specifically, the result that participants who opted to view LLM-generated fact checks were more likely to share all content and to believe false content was more pronounced among participants with positive attitudes towards AI. Similarly, participants who opted to view LLM fact checks about partisan false headlines were more likely to believe headlines that were congruent, as opposed to incongruent, with their partisan interests. For more information about these analyses, please see the Interaction analyses section in Supplementary information.

### v4

# Fact-checking information from large language models can decrease headline discernment

@@ Abstract
Fact checking can be an effective strategy against misinformation, but its implementation at scale is impeded by the overwhelming volume of information online. Recent artificial intelligence (AI) language models have shown impressive ability in fact-checking tasks, but how humans interact with fact-checking information provided by these models is unclear. Here, we investigate the impact of fact-checking information generated by a popular large language model (LLM) on belief in, and sharing intent of, political news headlines in a preregistered randomized control experiment. Although the LLM accurately identifies most false headlines (90%), we find that this information does not significantly improve participants’ ability to discern headline accuracy or share accurate news. In contrast, viewing human-generated fact checks enhances discernment in both cases. Subsequent analysis reveals that the AI fact-checker is harmful in specific cases: it decreases beliefs in true headlines that it mislabels as false and increases beliefs in false headlines that it is unsure about. On the positive side, AI fact-checking information increases the sharing intent for correctly labeled true headlines. When participants are given the option to view LLM fact checks and choose to do so, they are significantly more likely to share both true and false news but only more likely to believe false headlines. Our findings highlight an important source of potential harm stemming from AI applications and underscore the critical need for policies to prevent or mitigate such unintended consequences.
{adjustwidth}
.9cm.9cm
@@
This study explores how LLMs used for fact-checking affect the perception and dissemination of political news headlines. Despite the growing adoption of AI and tests of its ability to counter online misinformation, little is known about how people respond to LLM-driven fact-checking. This randomized control experiment reveals that even LLMs that accurately identify false headlines do not necessarily enhance users’ abilities to discern headline accuracy or promote accurate news sharing. LLM fact checks can actually reduce belief in true news wrongly labeled as false and increase belief in dubious headlines when the AI is unsure about an article’s veracity. These findings underscore the need for research on AI fact-checking’s unintended consequences, informing policies to enhance information integrity in the digital age.
@@
Digital misinformation has rapidly become a critical issue of modern society (1, 2). Recent work suggests that misinformation can erode support for climate change (3, 4), contribute to vaccine hesitancy (5, 6, 7), exacerbate political polarization (8), and even undermine democracy (9). As a mitigation strategy, fact checking has proved effective at reducing people’s belief in (10, 11, 12) and intention to share (13) misinformation in various cultural settings (14). However, this approach is not scalable, greatly limiting its applications (15).
@@
To tackle this challenge, researchers and social media platforms have been exploring automated methods (16) to detect misinformation (17, 18) and fact-check claims (19, 20, 16, 21, 22, 23). A robust fact-checking system must possess the ability to detect claims, retrieve relevant evidence, assess the veracity of each claim, and yield justifications for the provided conclusions (24, 25). Previous work attempting to meet these goals typically adopts cutting-edge artificial intelligence (AI) methods, specifically natural language processing. Nevertheless, building a functional system that can handle the vast volume of digital information on the internet, spanning various contexts and languages, remains a daunting task.
@@
Recent advances in large language models (LLMs) may appear to provide a feasible path forward. Trained on massive datasets of text from the internet, including news articles, books, and websites (26), these models are knowledgeable about a wide range of topics and have shown impressive performance on tasks such as text summarization and named entity recognition (27, 28). Outside the laboratory, LLMs have demonstrated remarkable abilities, even passing challenging exams designed for humans (29, 30).
@@
Analyses of ChatGPT, a prominent LLM, suggest it can rate the credibility of news outlets (31) and has great potential to fact-check claims (32, 33, 34), especially when augmented with additional data (35). Messages provided by LLMs to correct social media misinformation can be better than corrective messages generated by humans (36). These models can generate convincing justifications for the information they provide and even engage in conversations with users to provide additional context and facilitate understanding in multiple languages. Such capabilities of LLMs, coupled with open-sourcing efforts (37, 38), create a favorable environment for the development of scalable and reliable AI systems that can verify substantially more claims on the internet than is currently possible.
@@
However, realizing this potential requires humans to integrate LLMs into the digital information ecosystem effectively. Unfortunately, human-AI interaction is notoriously complex (39). Prior work has shown that AI is often seen as objective (40, 41, 42, 43), yet trust in AI depends on various factors such as individual expectations (44, 45), system interactivity (46, 47), and whether the AI provides information about its recommendations (48, 49).
@@
In the present context, it remains unclear how humans would interact with fact-checking information provided by state-of-the-art LLMs. Therefore, a thorough analysis of this misinformation intervention is necessary before deploying models in the wild. To this end, we conduct a preregistered (50), randomized controlled experiment to examine the causal effects of viewing fact-checking information provided by ChatGPT 3.5 on individual beliefs in and intention to share political news headlines. We selected ChatGPT for our study despite it not being specifically tailored for fact-checking. This decision was driven by its widespread public availability and use as well as the promising results emerging from tests of its claim verification capabilities at the time (34, 32, 33).
@@ Results
We recruited a representative sample of $N=2{,}159$ U.S. participants (see Methods for more information). All participants were presented with the same 40 real political news stories, which included a headline, lede sentence (if present), and image. Half of these headlines were true and the other half were false. Half were favorable to Democrats and the other half were favorable to Republicans (see Methods for details).
Participants were separated into “belief” and “sharing” groups in which they were asked to indicate, respectively, whether they believed headlines to be accurate or would be willing to share them on social media. The response options for both questions were “Yes” or “No.” These questions were asked separately as priming individuals to think about headline veracity can alter sharing behavior (51, 15). Each group included four conditions: a control group and three treatment conditions. In the human fact check condition, participants were presented with traditional fact checks generated by humans. The other two conditions emulated hypothetical scenarios for an automated fact-checking system on a social media platform: treated subjects were either forced to view fact-checking information provided by ChatGPT (LLM-forced) or given the option to reveal that information by clicking a button (LLM-optional). ChatGPT fact-checking information was identical for all treated subjects and presented directly below the corresponding headline. Participants in the LLM treatment conditions were informed that the fact checks were generated by ChatGPT, while those in the human fact checks condition were only informed that they would receive fact-checking information. Subjects in the control condition were only shown headlines and asked the belief/sharing question without being exposed to any fact-checking information. The experimental design is illustrated in Fig. 1a (see Methods for more details).
Unless otherwise stated, all $P$ values presented here are generated with two-tailed Mann-Whitney U tests and adjusted with Bonferroni correction for multiple comparisons. In the Supplementary Information, we also report on linear regression for all results, employing robust standard errors clustered on participant and headline.
@@
Although limited in size, our set of headlines provides us with a balanced representation of political biases and factual accuracy to evaluate the LLM. Overall, this analysis suggests that the LLM is an accurate fact-checker for false content. For true headlines, it is less accurate but can generally identify and explain when it cannot provide accurate fact-checking information. These results align with earlier studies that delve into the accuracy of LLM fact-checking utilizing much larger datasets (32, 33, 34).
An additional analysis of various prompt engineering methods and their performance, measured using standard binary classification metrics, is available in the Supplementary Information. Techniques that forced ChatGPT to produce only “true” or “false” judgments did not enhance the model’s overall accuracy.
@@
To evaluate the effectiveness of a misinformation intervention, it is crucial to measure its impact on belief in and sharing of both true and false headlines (52, 51). Although the veracity of headlines may not always fit neatly into true and false categories, as in the case of rumors with unclear veracity that are later clarified (53), this framework defines the desired outcome: an effective misinformation intervention should enhance individuals’ ability to distinguish between true and false headlines such that they believe/share more accurate news.
@@
To capture the causal effects of LLM-generated fact-checking information, we compare the average discernment of participants in the treatment conditions (LLM-forced and LLM-optional) to those in the control condition. Discernment is defined as the difference between the proportion of true and false headlines that participants believe (or are willing to share), capturing the intervention’s impact on both news categories. The inclusion of the human fact check condition allows us to differentiate AI-related effects from those associated with traditional fact checking.
@@
Figure 1 (panels c,d) illustrates the effects of fact checking on belief in and intent to share true and false headlines under each condition, including the mean group discernment as an annotation. In contrast with our preregistered expectations, discernment within both the belief and sharing groups was unaffected by the LLM treatment, regardless of condition. In the belief group (Fig. 1c), participants who were forced to view AI fact checks displayed a slight mean reduction ($-4.50$%) in discernment when compared to the control group ($U=31{,}993$, $P=0.61$, $d=-0.15$, 95% CI: $[-10.04\%,0.90\%]$). The discernment of those given the option to view fact checks in this group was virtually unaffected, decreasing on average by only $-0.27$% ($U=31{,}265$, $P=1$, $d=-0.01$, 95% CI: $[-5.52\%,5.03\%]$).
@@
We observe similar results regarding sharing behavior. Participants in the LLM-forced and LLM-optional conditions of the sharing group displayed a mean reduction of $-0.43$% ($U=35{,}785$, $P=1$, $d=-0.02$, 95% CI: $[-3.93\%,3.05\%]$) and $-0.67$% ($U=34{,}860$, $P=1$, $d=-0.03$, 95% CI: $[-4.18\%,2.92\%]$), respectively, when compared with the control group.
In contrast to the above, we observe a significant increase in discernment for participants who viewed human fact checks within both the belief and sharing groups. On average, belief discernment increased by 18.06% ($U=25{,}210$, $P<0.001$, $d=0.50$, 95% CI: $[12.00\%,24.00\%]$), and sharing discernment increased by 8.98% ($U=34{,}224$, $P=0.001$, $d=0.35$, 95% CI: $[4.89\%,13.33\%]$). These effects are primarily due to an increased belief in and willingness to share true headlines—which increased by 15.63% and 9.58%, respectively—while the impact on false headlines remained largely unchanged.
@@
In summary, these results indicate that human fact checks served as an effective misinformation intervention, while those generated by the LLM did not. This is unexpected, considering that the AI provides participants with useful information, particularly for false headlines. However, this analysis does not account for the accuracy of the AI’s responses, nor does it examine how behaviors vary when individuals choose to view or not view this information. To delve deeper into these dynamics, we have supplemented our preregistered design with two additional exploratory analyses in the sections that follow.
@@
Here, we explore the causal effects of viewing LLM fact-checking information when accounting for model accuracy. The judgments made by ChatGPT for both true and false headlines fall into one of three categories: correct, incorrect, or unsure. This results in six different scenarios (True/False $\times$ Correct/Incorrect/Unsure) in which effects may be observed. However, our data contain no false headlines judged by ChatGPT to be true, resulting in five scenarios for each previously considered comparison (Belief/Sharing $\times$ Control/LLM-optional/LLM-forced).
To evaluate the potential impact of LLM-generated fact checks, we compare the LLM-forced and control conditions in these five scenarios, as illustrated in Fig. 2. Annotations indicate mean group differences and highlight the significant effects identified through Bonferroni-adjusted Mann-Whitney U tests.
@@
In the belief group, we found significant undesirable effects showing that LLM fact checks decreased participants’ discernment. Specifically, there was a 12.75% decrease in the belief of true headlines incorrectly judged as false by ChatGPT ($U=35{,}937$, $P<0.001$, $d=-0.38$, 95% CI: $[-18.67\%,-6.89\%]$) and a 9.12% increase in the belief of false headlines where the AI expressed uncertainty ($U=25{,}931$, $P=0.03$, $d=0.22$, 95% CI: $[1.69\%,16.35\%]$). Both cases demonstrate a behavioral change that is counter to the ideal outcomes of any misinformation intervention.
Regarding the sharing group, we observed mixed results. While there was an 11.09% increase in participants’ intention to share true headlines correctly judged by ChatGPT ($U=30{,}897$, $P=0.02$, $d=0.26$, 95% CI: $[4.02\%,18.06\%]$), there was also a 9.77% increase in the intention to share false headlines where ChatGPT expressed uncertainty ($U=31{,}856$, $P=0.05$, $d=0.22$, 95% CI: $[2.31\%,17.25\%]$). The former increases sharing discernment, while the latter reduces it by a similar amount.
@@
We next analyze participants’ behavior in the LLM-optional conditions, comparing those who opt in to see LLM fact-checking information versus those who opt out.
On average, participants chose to view fact checks for slightly more than half of the headlines. The mean number of fact checks viewed for the belief and sharing groups was 21.6 (SD = 15.8) and 23.8 (SD = 15.7), respectively. However, the distribution was bimodal: about half the participants viewed fact checks for most headlines, while the other half viewed them for only a handful. Participants who viewed fact checks for more than half of the 40 headlines (52.1%) averaged viewing 36.7 (SD = 5.5). In contrast, those who viewed less than half averaged 7.5 views (SD = 6.4). A Mann-Whitney U test revealed no significant difference in opt-in behavior between true and false headlines ($P=0.1$). See the Supplementary Information for more details.
Figure 3 illustrates the belief in and intention to share headlines for which subjects chose to see versus not see LLM fact checks, for both true and false headlines. Each subject’s contribution to the group mean values and confidence intervals are weighted by the number of times they chose to see (or not see) each type of headline. Figure 3a shows that participants who chose to see LLM fact checks were significantly more likely to believe false headlines accurately identified by the LLM, with a 29.35% increase ($U=23{,}480$, $P=0.005$, $d=0.63$, 95% CI: $[20.81\%,37.93\%]$), as well as those the model was unsure about, with a 28.12% increase ($U=14{,}260$, $P<0.001$, $d=0.64$, 95% CI: $[18.43\%,38.12\%]$). There was no significant difference in belief for true headlines that the model could not classify, with a 5.51% increase ($U=20{,}452$, $P=1$, $d=0.12$, 95% CI: $[-3.70\%,14.47\%]$), and we only observe a 7.46% increase for accurately identified true headlines ($U=10{,}941$, $P=0.04$, $d=0.18$, 95% CI: $[-1.50\%,16.35\%]$). However, a significant decrease for those misjudged as false (16.59% decrease; $U=10{,}299$, $P<0.001$, $d=-0.35$, 95% CI: $[-26.34\%,-7.24\%]$) was found. Figure 3b shows that participants who chose to see LLM fact checks were significantly more likely to share headlines in all scenarios. These increases ranged from 29% for true headlines judged as false to 39% for true headlines judged as true (see the Supplementary Information for statistics).
Figure 3: Proportions of headlines that participants in the optional condition indicated they (a) believed or (b) were willing to share on social media. These proportions are based on the headline’s veracity, whether participants chose to see LLM fact-checking information (opt in) or not (opt out), and how the LLM judged the headlines (True, Unsure, False). No false headlines were judged as true. Error bars represent 95% confidence intervals, calculated using a weighted bootstrapping method with 5,000 resamples. The mean difference between opt-in and opt-out groups (rounded to whole percentage) is annotated for each condition.
We note that this particular within-group analysis does not allow us to identify causal effects because participants are not randomly assigned to the treatment (opt in) or comparison (opt out) group for each headline. Nonetheless, when participants viewed LLM-generated fact-checking information, they were more likely to share both true and false news. Additionally, those who viewed this information were less likely to believe true news misjudged as false and more likely to believe false news, even when accurately identified as such by the model.
@@
Our preregistered analyses also examined the potential roles of individual attitudes toward AI (ATAI) and the partisan congruence of headlines. While we find minimal evidence that these variables significantly impacted the results of our first two analyses, we observed specific relationships when individuals had the option to view LLM fact-checking information.
In particular, we found clear evidence that participants with positive attitudes toward AI who chose to view LLM-generated fact checks were significantly more likely to share those headlines across all fact-checking scenarios. However, the relationship between ATAI and participant belief was less clear. Nonetheless, the tendency for participants to share and believe true news that the LLM was unsure about was more pronounced among those with positive attitudes towards AI when viewing AI fact checks.
When participants encountered politically incongruent true headlines that the LLM was unsure about, their likelihood of believing or sharing them diminished significantly. This relationship persisted irrespective of whether participants opted to access the fact-checking information. We observed a similar negative relationship in only one other scenario: for incongruent false headlines when participants did not view LLM fact checks. For more details on these analyses, please refer to the Supplementary Information.